The acceleration of a motorcycle is given by  ax(t)=At-Bt2, where A=1.50m/s3  and B=0.12m/s4 . The motorcycle is at rest at the origin at time t=0.

a. Find its position and velocity as a function of time.

b. Calculate the maximum Velocity

 


1
Expert's answer
2021-02-15T00:54:55-0500

(a) We can find the velocity of the motorcycle by taking the integral of the acceleration over time:


"v(t)=\int a(t)dt,""v(t)=\int (At-Bt^2)dt=\dfrac{At^2}{2}-\dfrac{Bt^3}{3}."

We can find the position of the motorcycle by taking the integral of the velocity over time:


"s(t)=\int v(t)dt,""s(t)=\int (\dfrac{At^2}{2}-\dfrac{Bt^3}{3})dt=\dfrac{At^3}{6}-\dfrac{Bt^4}{12}."

(b) Let's first take the derivative from the velocity with respect to time:


"\dfrac{d}{dt}(\dfrac{At^2}{2}-\dfrac{Bt^3}{3})=At-Bt^2."

Let's set the derivative equal to zero and find the time at which the velocity of the motorcycle is maximum:


"At-Bt^2=0,""t(A-Bt)=0."

This equation has two roots: "t=0" and "t=\dfrac{A}{B}." Since, at "t=0" the velocity of the motorcycle is zero, its velocity is maximum at "t=\dfrac{A}{B}".

Substituting "t" into the equation for "v(t)" we get:


"v_{max}=\dfrac{A(\dfrac{A}{B})^2}{2}-\dfrac{B(\dfrac{A}{B})^3}{3}=\dfrac{A^3}{6B^2}."

Let's substitute the numbers:


"v_{max}=\dfrac{(1.50\ \dfrac{m}{s^3})^3}{6\cdot(0.12\ \dfrac{m}{s^4})^2}=39.1\ \dfrac{m}{s}."